An oxygen sample occupies 4.75L at 25 Celsius. Predict the volume this sample will occupy at 75 Celsius assuming the pressure stays constant.
Answer #1

At constant pressure ; volume is directly proportional to temperature.

V~ T

i.e

V1/T1 = V2/T2

V1 = 4.75 L

T1 = 25 + 273 K = 298K

V2 = ?

T2 = 75 + 273K = 348K

4.75 / 298 = V2 / 348

V2 = 5.55 L